40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6533     Entry time: Fri Apr 13 01:27:10 2012
Author: Jenne 
Type: Update 
Category: SUS 
Subject: Oplevs recentered 

It's been a while since I think anyone has done it, and several optics were pretty far from centered, so I centered all of the oplevs except for SRM.

I am confident about my arm alignment, MICH and PRC (so BS, ITMX, ITMY, PRM, ETMX and ETMY), but I wasn't sure if I was getting SRC right, so I didn't touch the SRM's oplev.

Suresh removed all of the IPPOS measurement optics, so there was nothing blocking the BS and PRM oplevs.

However, the PRM oplev was ridiculously bad, and I don't know how long it's been that way.  Some of the optics shooting the beam into the chamber weren't optimally aligned, so the beam coming out of the chamber was hitting the lowest edge of the optic mount, for the first optic the beam encountered.  I adjusted the mirror launching the beam into the chamber by a teeny bit, so that the outcoming beam was ~horizontal and hitting the center of the first steering mirror in pitch.  I had to move that steering mirror a little to the right (if you are staring at the HR face of the mirror), to get the beam to come close to the horizontal center of the optic.  Then I proceeded to do normal oplev alignment.

Also, I've noticed lately that ITMX is noisier than all the other optics.  It's kind of annoying.  The sensor RMS values reported for the ITMX watchdogs for UL and LL are rarely below 2, and are often (~70% of the time?) above 3.  The SD RMS is a normal 1-ish.

ELOG V3.1.3-